OpenStudy (anonymous):

Solve for roots using theory of equations- x^4 - 16x^3 + 86x^2 - 176x + 105 = 0 , where two roots are 1 and 7.Find the others.

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

since you already know two roots, you can reduce that to a quadratic, correct?

OpenStudy (anonymous):

but i want by the method of sum of roots and product of roots.

OpenStudy (anonymous):

the product of the roots is 105... but i'm not familiar with the "sum of roots" for a fourth degree polynomial... sorry..

OpenStudy (anonymous):

i think u can divide the 4th degree polynomial by (x-1)(x-7) and it reduce to a 2nd degree polynomial .

OpenStudy (anonymous):

ok... i googled it.... 7 * 1 * r1 * r2 = 105 7 + 1 + r1 + r2 = 16 <--- b/a two equations, two unknowns... solve the system for r1 and r2....

OpenStudy (anonymous):

so r1 * r2 = 15 r1 + r2 = 8 can you take it from here?

OpenStudy (anonymous):

ok